# T-Explainer: A Model-Agnostic Explainability Framework Based on Gradients

### v1

@@ Abstract
The development of machine learning applications has increased significantly in recent years, motivated by the remarkable ability of learning-powered systems to discover and generalize intricate patterns hidden in massive datasets. Modern learning models, while powerful, often exhibit a level of complexity that renders them opaque black boxes, resulting in a notable lack of transparency that hinders our ability to decipher their decision-making processes. Opacity challenges the interpretability and practical application of machine learning, especially in critical domains where understanding the underlying reasons is essential for informed decision-making. Explainable Artificial Intelligence (XAI) rises to meet that challenge, unraveling the complexity of black boxes by providing elucidating explanations. Among the various XAI approaches, feature attribution/importance XAI stands out for its capacity to delineate the significance of input features in the prediction process. However, most existing attribution methods have limitations, such as instability, when divergent explanations may result from similar or even the same instance. In this work, we introduce T-Explainer, a novel local additive attribution explainer based on Taylor expansion endowed with desirable properties, such as local accuracy and consistency, while stable over multiple runs. We demonstrate T-Explainer’s effectiveness through benchmark experiments with well-known attribution methods. In addition, T-Explainer is developed as a comprehensive XAI framework comprising quantitative metrics to assess and visualize attribution explanations.
Keywords: Black-box models, explainable artificial intelligence, XAI, interpretability, local explanations.This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ 1 Introduction
Artificial Intelligence is not a vision for the future. It is our present reality. Terms such as Neural Networks, Machine Learning, Deep Learning, and other facets of the Artificial Intelligence universe have seamlessly shifted from futuristic concepts to near-ubiquitous elements in our daily discourse. The confluence of recent strides in hardware processing capabilities, abundant data accessibility, and refined optimization algorithms has facilitated the creation of intricate and non-linear machine learning models [15]. In the contemporary landscape, those models have achieved unprecedented performance levels, surpassing what was deemed inconceivable just a few years ago, outperforming human abilities and previously known methods in central research areas [67, 63, 56].
@@
Widely used machine learning models are impenetrable as far as simple interpretations of their mechanisms go [16]. In this context, Explainable Artificial Intelligence (XAI) has emerged to address the challenges outlined above, seeking to provide human-understandable insights into the complexities of models that are inherently challenging to interpret. Significantly, the field of explainability has progressed with the introduction of innovative methodologies, and research focused on discerning the strengths and limitations of XAI models [2]. Feature attribution/importance methods [60] are of particular relevance, aiming to quantify the contribution of individual input variables on predictions made by black-box models [2].
While feature attribution methods prove valuable, they come with drawbacks that diminish the trust and confidence placed in their application. For instance, different methods can result in distinct explanations for the same data instances, making it difficult to decide which outcome to believe [33, 60]. Moreover, many well-known methods, including SHAP [40] and LIME [48], suffer from instability, producing different explanations from different runs on a fixed machine learning model and dataset [23].
@@
In this work, we introduce T-Explainer, a novel post-hoc explanation method that relies on the solid mathematical foundations of Taylor expansion to perform local and model-agnostic feature importance attributions. T-Explainer is a local additive explanation technique with desirable properties such as local accuracy, missingness, and consistency [40]. Furthermore, T-Explainer is entirely deterministic, guaranteeing stable results across multiple runs and delivering consistent explanations for similar instances.
To evaluate the quality and usefulness of the T-Explainer, we performed several quantitative comparisons against well-known local feature attribution methods, assessing the stability and additivity preservation of such XAI methods. We implemented the T-Explainer integrated with the visualization tools available in the SHAP library, incorporating a suite of quantitative metrics to evaluate feature importance explanations, rendering the T-Explainer a comprehensive XAI framework. In summary, the main contributions of this work are:
@@
  • A comprehensive set of comparisons against well-known local feature attribution methods using quantitative metrics, aiming mainly to assess stability and additivity preservation.
  • A Python library that integrates T-Explainer with other explanation tools, making the proposed method easy to use in different applications. The framework, datasets, and all related materials are openly available online111 The GitHub link will be provided upon acceptance.
@@ 2 Related work
Several XAI techniques have been proposed to deal with black-box models. To contextualize our contributions, we focus the following discussion on post-hoc feature importance/attribution methods [18, 68, 33]. A more comprehensive discussion about XAI can be found in several surveys summarizing existing approaches and their properties [2, 7, 38, 12, 17] and metrics for evaluating explanation methods [70, 43, 65, 4].
@@
SHAP has a solid theoretical foundation derived from game theory [52] that grants SHAP desirable properties such as local accuracy, missingness, and consistency [40]. However, the exact computation of SHAP values is NP-hard, demanding Monte Carlo sampling-based approximations [60], which introduces instability similar to the LIME method, even in its model-specific variants [41, 19, 20]. To avoid instability, deterministic versions of LIME [71], optimization [35], and learning-based [55] approaches have been proposed, but with the price of increasing the number of parameters to be tuned.
@@ 3 The T-Explainer
In this Section, we introduce the theoretical foundations, properties, and computational aspects of the proposed T-Explainer method.
Let $\mathbf{X}$ be a multidimensional dataset where each data instance $\mathbf{x}=(x_{1},\dots,x_{n})\in\mathbf{X}$ is a vector in $\mathbb{R}^{n}$ and $f$ be a machine learning model. For simplicity, let’s assume that $f$ is a binary classification model trained on $\mathbf{X}$, that is, $f(\mathbf{x})\in[0,1]$ accounts for the probability of $\mathbf{x}$ belonging to class $1$, and $(1-f(\mathbf{x}))$ is the probability of belonging to class 00. All the following reasoning can be extended to regression models.
@@
$f:\mathbf{X}\rightarrow[0,1]\subset{\rm I\!R}$ (1)
@@
The $i$-th gradient element corresponds to the partial derivative of $f$ concerning the $i$-th attribute of $\mathbf{x}$. Note that the gradient of $f$ at $\mathbf{x}$ corresponds to the Jacobian matrix when $f$ is a real-valued function. The dot product between the gradient and the displacement vector $\mathbf{h}$ is a linear map from $\mathbb{R}^{n}$ to $\mathbb{R}$, being well-known as the best linear approximation of $f$ in a small neighborhood of $\mathbf{x}$ [44, 45]. Therefore, the gradient can be used to analyze how small perturbations in the input data affect the model output. The right side of Equation 2 is a linear equation that approximates the behavior of $f$ nearby $\mathbf{x}$, and by being a linear mapping, it is naturally interpretable. In other words, the gradient of a model $f$ can be used to generate explanations.
The formulation above resembles the Vanilla Gradient [54] Taylor expansion-based procedure to compute saliency maps. The difference is that, in our case, the attributions are not dependent on class information and do not rely on further parameters specific to the model’s architecture. In addition, T-Explainer differs from previous gradient-based methods by relying on additive explanation modeling and a deterministic optimization procedure to approximate gradients, as described below.
Let $\mathbf{h}=\mathbf{z}^{\prime}\in\mathbb{R}^{n}$ be a perturbation in $\mathbf{x}$, that is, $\mathbf{x}^{\prime}=\mathbf{x}+\mathbf{z}^{\prime}$ is a point in a small neighborhood of $\mathbf{x}$. The T-Explainer can thus be defined as an additive explanation modeling $g_{\mathbf{x}}$, given by:
@@
where $\phi_{0}=\mathrm{\mathbb{E}}[f(\mathbf{X})]$ represents the expected prediction value, and $\phi_{i}=\frac{\partial{f}(\mathbf{x})}{\partial x_{i}}$. The prediction expected value is a statistical value that is not trivial to estimate for an arbitrary dataset. In practice, it is computed through the average model output across the training dataset $\mathbf{X}$ when the feature values $\mathbf{X}_{i}$ are unknown. As a fundamental property of additive explanations, Equation 4 approximates the original predicted value $f(\mathbf{x})$ locally by summing its feature importances [40].
@@ 3.2 T-Explainer: Computational Aspects
$FD_{f(\mathbf{x})}=\frac{f(\mathbf{x}+\mathbf{h})-f(\mathbf{x}-\mathbf{h})}{2\mathbf{h}}.$ (9)
@@
Specifically, we rely on the centered finite difference because that approach simply averages two one-sided perturbations of each attribute, resulting in a second-order accurate approximation with an error proportional to $\mathbf{h}^{2}$ [37]. The centered finite difference approximation of the partial derivatives demands to set up the magnitude of the perturbation $\mathbf{h}$. The displacement $\mathbf{h}$ must be small enough to generate a perturbation close to the input data $\mathbf{x}$. By definition, the derivatives of $f$ are computed by making $\mathbf{h}\to 0$ in Equation 9. In other words, we have to set $\mathbf{h}$ to a small value.
@@
where $\mathcal{L}$ is the cost function (MSE in our case), $\nabla f_{\mathbf{x}}$ represents the approximated value of the gradient $\nabla f(\mathbf{x})$ in each iteration of the $\mathbf{h}$ optimizer algorithm computing the method in Equation 9 on Equation 2, $\epsilon_{\mathbf{x}}$ is the cost threshold upper-bounded by $h_{min}^{2}$ [37], and $\theta$ is a method that ensures the numerical stability of the optimization process by checking if $\mathbf{h}$ is keeping the gradient (Jacobian matrix) $\nabla f_{\mathbf{x}}$ non-singular (or full-rank). It is important to highlight that local perturbations must preserve the data’s normalization range to avoid model extrapolations [31]. Further details about the optimization process are given in the Appendix Section. This approach can find the gradient ${\nabla f(\mathbf{x})}$ of a binary-class model or even the entire Jacobian related to all the model’s output probabilities. Once the gradient is estimated in $\mathbf{x}$, the importance of each attribute is computed as $\phi_{i}=\frac{\partial{f}(\mathbf{x})}{\partial x_{i}}$.
@@ 4 Quantitative Metrics
In this Section, we describe the evaluation metrics integrated into the T-Explainer framework. Relative Input Stability (RIS) and Relative Output Stability (ROS) [3, 4] are metrics used to evaluate the stability of local explanations to changes in input data and output prediction probabilities, respectively. We followed the basic definitions in Agarwal et al. [3] but introduced improvements to address the issues described below.
Given $\mathcal{N}_{\mathbf{x}}$ a neighborhood of perturbed instances $\mathbf{x}^{\prime}$ around $\mathbf{x}$, with $e_{\mathbf{x}}$ and $e_{\mathbf{x}^{\prime}}$ representing the attribution vectors explaining $\mathbf{x}$ and $\mathbf{x}^{\prime}$, and $f({\mathbf{x}})$ and $f({\mathbf{x}^{\prime}})$ the output prediction probabilities of $\mathbf{x}$ and $\mathbf{x}^{\prime}$, RIS and ROS metrics are formally defined as:
@@
$\forall\mathbf{x}^{\prime}\in\mathcal{N}_{\mathbf{x}}$, with ${p}$ defining the $l_{p}$ norm used to measure the changes and $\epsilon_{c}>0$ as a clipping threshold to avoid zero division. The larger the RIS or ROS values, the more unstable the method is related to input or output prediction perturbation.
@@
RIS and ROS handle values close to zero in the normalization processes. However, the original metrics implemented in Agarwal et al. [4] applied a clip method that ignores significant negative values by handling any negative value as a small value. Then, we designed a clip method that preserves significant values, both positive and negative, while discarding those close to zero. The original implementations may also ignore instances near the model’s decision boundaries. We solved that issue by sampling perturbations of each $\mathbf{x}$, sorting the perturbed data based on the distance to $\mathbf{x}$, and generating a neighborhood holding a minimum set of perturbed instances. This way, we ensure coverage and avoid measurement omissions or potential crashes in the metrics. Additionally, we extended RIS and ROS to evaluate the mean changes and the respective standard deviations, in addition to the original maximum change in explanations.
A novel metric, Run Explanation Stability (RES), was added to our framework to measure the local stability through multiple runs on non-perturbed inputs. RES assesses the consistency of several explanations for the same instance under the same settings, with higher values indicating lower stability rates. Given $n$ explanations $g_{\mathbf{x}i},\dots,g_{\mathbf{x}n}$ of the same instance $\mathbf{x}$, with each $g_{\mathbf{x}i}$ as a scalar representing the prediction approximation (see Equation 4), let $\bar{g}_{\mathbf{x}}$ be the mean explanation from those $n$ explanations. We define the RES metric as:
$\text{RES}(g_{\mathbf{x}},\bar{g}_{\mathbf{x}})=\max_{k=1}^{n}{\|\bar{g}_{\mathbf{x}}-g_{\mathbf{x}k}\|}\text{,}\>\>\>\forall\mathbf{x}\in\mathbf{X}\text{.}$ (13)
@@
The Reiteration Similarity metric inspired the rationale behind RES [6]; however, we simplified the formulation using standard deviation as the similarity measure.
Another new addition is the Faithfulness of an Additive Explanator (FAE) metric, which assesses how an additive feature importance explainer preserves the primary property of local accuracy, i.e., the model prediction should be reconstructed by the summation of the importance values as in Equation Equation 5 [40]. Given an output prediction probability $f(\mathbf{x})$ and its explanation $g_{\mathbf{x}}$, we define the FAE metric as the frequency or ratio in which $\|f(\mathbf{x})-g_{\mathbf{x}}\|\geq\epsilon$, $\forall\mathbf{x}\in\mathbf{X}$, with $\epsilon$ a tolerance threshold. Aiming to analyze better each model’s additiveness, the FAE metric also computes the mean values in which $\|f(\mathbf{x})-g_{\mathbf{x}}\|$. In this case, the smaller the FAE mean, the closer the approximations between $f(\mathbf{x})$ and $g_{\mathbf{x}}$ will be from the threshold $\epsilon$ and, thus, the more “additive” the model is (the explainer is more faithful regarding the local accuracy property).
@@
RIS, ROS, RES, and FAE metrics are unitless quantifiers with no “ideal” desirable values. Reasonable values depend on the context of the application, meaning those metrics must be interpreted relatively by comparing their results across XAI methods [5].

### v2

@@ Abstract
The development of machine learning applications has increased significantly in recent years, motivated by the remarkable ability of learning-powered systems to discover and generalize intricate patterns hidden in massive datasets. Modern learning models, while powerful, often have a level of complexity that renders them opaque black boxes, resulting in a notable lack of transparency that hinders our ability to decipher their reasoning. Opacity challenges the interpretability and practical application of machine learning, especially in critical domains where understanding the underlying reasons is essential for informed decision-making. Explainable Artificial Intelligence (XAI) rises to address that challenge, unraveling the complexity of black boxes by providing elucidating explanations. Among the various XAI approaches, feature attribution/importance stands out for its capacity to delineate the significance of input features in the prediction process. However, most existing attribution methods have limitations, such as instability, when divergent explanations may result from similar or even the same instance. This work introduces T-Explainer, a novel local additive attribution explainer based on Taylor expansion. It has desirable properties, such as local accuracy and consistency, making T-Explainer stable over multiple runs. We demonstrate T-Explainer’s effectiveness in quantitative benchmark experiments against well-known attribution methods. Additionally, we provide several tools to evaluate and visualize explanations, turning T-Explainer into a comprehensive XAI framework.
Keywords: Black-box models, explainable artificial intelligence, XAI, interpretability, local explanations.
@@ 1 Introduction
Artificial Intelligence is not a vision for the future. It is our present reality. Terms such as Neural Networks, Machine Learning, Deep Learning, and other facets of the Artificial Intelligence universe have seamlessly shifted from futuristic concepts to near-ubiquitous elements in our daily discourse. The confluence of recent strides in hardware processing capabilities, abundant data accessibility, and refined optimization algorithms has facilitated the creation of intricate and non-linear machine learning models [16]. In the contemporary landscape, those models have achieved unprecedented performance levels, surpassing what was deemed inconceivable just a few years ago, outperforming human abilities and previously known methods in central research areas [63, 56].
@@
Widely used machine learning models are impenetrable as far as simple interpretations of their mechanisms go [17]. In this context, Explainable Artificial Intelligence (XAI) has emerged to address the challenges outlined above, seeking to provide human-understandable insights into the complexities of models that are inherently challenging to interpret. Significantly, the field of explainability has progressed with the introduction of innovative methodologies, and research focused on discerning the strengths and limitations of XAI models [2]. Feature attribution/importance methods are particularly relevant, as they are currently the most common explanation type [61, 42]. Such methods aim to quantify the contribution of individual input variables to predictions made by black-box models, providing local insights about the model reasoning [43].
While feature attribution methods prove valuable, they come with drawbacks that diminish the trust and confidence placed in their application. For instance, different methods can result in distinct explanations for the same data instances, making it difficult to decide which outcome to believe [61]. Moreover, many well-known methods, including SHAP [39] and LIME [47], suffer from instability, producing different explanations from different runs on a fixed machine learning model and dataset [22].
@@
In this work, we introduce T-Explainer, a novel post-hoc explanation method that relies on the solid mathematical foundations of Taylor expansion to perform local and model-agnostic feature importance attributions. T-Explainer is a local additive explanation technique with desirable properties such as local accuracy, missingness, and consistency [39]. Furthermore, T-Explainer is deterministic, guaranteeing stable results across multiple runs and delivering consistent explanations for similar instances.
We conducted several quantitative comparisons against well-known feature attribution methods using controlled synthetic and real-world datasets to evaluate the quality and usefulness of T-Explainer explanations. Our evaluations focused on the consistency, continuity, and correctness properties of the explanations’ content [42]. Since T-Explainer is an additive method [39] such as SHAP, we also performed an additivity preservation comparison. As a result of such an extensive evaluation process, we implemented a suite of quantitative metrics to assess different dimensions of feature importance explainers. These metrics were incorporated into the T-Explainer package, rendering it not just another XAI method but a comprehensive XAI framework.
In summary, the main contributions of this work are:
@@
  • A comprehensive set of comparisons against well-known local feature attribution methods using multiple quantitative metrics, aiming to assess stability, faithfulness, and additivity preservation.
  • A Python framework that integrates T-Explainer with other explainability tools, making the proposed method easy to use and assess in different applications. The T-Explainer, evaluation metrics, datasets, and all related materials are openly available online111 A GitHub link will be provided upon acceptance. .
@@ 2 Related work
Several XAI techniques have been proposed to deal with black-box models, either locally or globally [61]. To contextualize our contributions, we focus the following discussion on post-hoc feature importance/attribution methods [66, 32]. A more comprehensive discussion about XAI can be found in several surveys summarizing existing approaches and their properties [2, 8, 37, 43] and metrics for evaluating explanation methods [67, 4, 42].
@@
SHAP has a solid theoretical foundation derived from game theory [51] that grants SHAP desirable properties such as local accuracy, missingness, and consistency [39]. However, the exact computation of SHAP values is NP-hard, demanding Monte Carlo sampling-based approximations [61, 43], which introduces instability similar to the LIME method, even in its model-specific variants [40, 18, 19]. To avoid instability, deterministic versions of LIME [68], optimization [34], and learning-based [55] approaches have been proposed, but with the price of increasing the number of parameters to be tuned.
@@ 3 The T-Explainer
In this Section, we introduce the theoretical foundations, properties, and computational aspects of the T-Explainer method.
Let $\mathbf{X}$ be a multidimensional dataset where each data instance $\mathbf{x}=(x_{1},\dots,x_{n})\in\mathbf{X}$ is a vector in $\mathbb{R}^{n}$ and $f$ be a machine learning model. For simplicity, let’s assume that $f$ is a binary classification model trained on $\mathbf{X}$, that is, $f(\mathbf{x})\in(0,1)$ accounts for the probability of $\mathbf{x}$ belonging to class $1$, and $(1-f(\mathbf{x}))$ is the probability of belonging to class 00. All the following reasoning can be extended to regression models.
@@
$f:\mathbf{X}\rightarrow(0,1)\subset{\rm I\!R}$ (1)
@@
The $i$-th gradient element corresponds to the partial derivative of $f$ concerning the $i$-th attribute of $\mathbf{x}$. Note that the gradient of $f$ at $\mathbf{x}$ corresponds to the Jacobian matrix when $f$ is a real-valued function. The dot product between the gradient and the displacement vector $\mathbf{h}$ is a linear map from $\mathbb{R}^{n}$ to $\mathbb{R}$, being well-known as the best linear approximation of $f$ in a small neighborhood of $\mathbf{x}$ [45]. Therefore, the gradient can be used to analyze how small perturbations in the input data affect the model output. The right side of Equation 2 is a linear equation that approximates the behavior of $f$ nearby $\mathbf{x}$, and by being a linear mapping, it is naturally interpretable. In other words, the gradient of a model $f$ can be used to generate explanations.
The formulation above resembles the Vanilla Gradient [54] Taylor expansion-based procedure to compute saliency maps. The difference is that, in our case, the attributions are not dependent on class information and do not rely on further parameters specific to the model’s architecture. In addition, T-Explainer differs from previous gradient-based methods by relying on additive modeling and a deterministic optimization procedure to approximate gradients, as detailed below.
Let $\mathbf{h}=\mathbf{z}^{\prime}\in\mathbb{R}^{n}$ be a perturbation in $\mathbf{x}$, that is, $\mathbf{x}^{\prime}=\mathbf{x}+\mathbf{z}^{\prime}$ is a point in a small neighborhood of $\mathbf{x}$. The T-Explainer is defined as an additive explanation modeling $g_{\mathbf{x}}$, given by:
@@
where $\phi_{0}=\mathrm{\mathbb{E}}[f(\mathbf{X})]$ represents the expected prediction value, and $\phi_{i}=\frac{\partial{f}(\mathbf{x})}{\partial x_{i}}$. The prediction expected value is a statistical value that is not trivial to estimate for an arbitrary dataset. In practice, it is computed through the average model output across the training dataset $\mathbf{X}$ when the feature values $\mathbf{X}_{i}$ are unknown. Specifically, it does not mean that $\mathbf{X}_{i}=0$; it means we do not know the value of $\mathbf{X}_{i}$. Therefore, its distribution is estimated based on the data by taking the mean value and then averaging the predicted probabilities for each label. As a fundamental property of additive explanations, Equation 4 approximates the original predicted value $f(\mathbf{x})$ locally by summing its feature importances [39].
@@ 3.2 T-Explainer: Computational Aspects
$FD_{f(\mathbf{x})}=\frac{f(\mathbf{x}+\mathbf{h})-f(\mathbf{x}-\mathbf{h})}{2||\mathbf{h}||}.$ (9)
@@
Specifically, we rely on the centered finite difference because that approach simply averages two one-sided perturbations of each attribute, resulting in a second-order accurate approximation with an error proportional to $||\mathbf{h}||^{2}$ [36]. The centered finite difference approximation of the partial derivatives demands to set up the magnitude of the perturbation $\mathbf{h}$. The displacement $\mathbf{h}$ must be small enough to generate a perturbation close to the input data $\mathbf{x}$. By definition, the derivatives of $f$ are computed by making $\mathbf{h}\to 0$ in Equation 9. In other words, we have to set $\mathbf{h}$ to a small value.
@@
where $\mathcal{L}$ is the cost function (MSE in our case), $\nabla f_{\mathbf{x}}$ represents the approximated value of the gradient $\nabla f(\mathbf{x})$ in each iteration of the $\mathbf{h}$ optimizer algorithm computing the method in Equation 9 on Equation 2, $\epsilon_{\mathbf{x}}$ is the cost threshold upper-bounded by $(h_{min})^{2}$ [36], and $\theta$ is a method that ensures the numerical stability of the optimization process by checking if $\mathbf{h}$ is keeping the gradient (Jacobian matrix) $\nabla f_{\mathbf{x}}$ non-singular (or full-rank). It is important to highlight that local perturbations must preserve the data’s normalization range to avoid model extrapolations [30]. Further details about the optimization procedure are given in the Appendix Section. This approach can find the gradient ${\nabla f(\mathbf{x})}$ of a binary-class model or even the entire Jacobian related to all the model’s output probabilities. Once the gradient is estimated in $\mathbf{x}$, the importance of each attribute is computed as $\phi_{i}=\frac{\partial{f}(\mathbf{x})}{\partial x_{i}}$.
@@ 4 Quantitative Metrics
A common drawback of XAI is the lack of quantitative evaluation of XAI methods. Most proposals rely on visual inspections or simplified human-centered case studies to check whether an explanation “looks reasonable.” However, such anecdotal strategies are prone to subjectivities and do not allow for formal comparisons between explanation methods, which is insufficient to ensure a robust verification of explanations’ consistency, continuity, and correctness [42]. This Section describes the evaluation metrics integrated into the T-Explainer framework. We focused our evaluations on automated quantitative metrics that enable us to compare different explanation methods regarding their stability, faithfulness, and local accuracy preservation.
A stability metric measures the robustness or sensitivity of an explainer when exposed to slightly different versions of an original input sample. Relative Input Stability (RIS) and Relative Output Stability (ROS) [3, 4] are metrics used to evaluate the stability of local explanations to changes in input data and output prediction probabilities, respectively. Alvarez-Melis and Jaakkola [6] introduced the stability concept, which Agarwal et al. [3] recently enhanced through the “relative stability” formulation. We followed theoretical definitions of RIS and ROS as defined by Agarwal et al. [4]; however, we introduced improvements to address some identified issues, which are detailed below.
Given $\mathcal{N}_{\mathbf{x}}$ a neighborhood of perturbed samples $\mathbf{x}^{\prime}$ around an original input instance $\mathbf{x}$, with $e_{\mathbf{x}}$ and $e_{\mathbf{x}^{\prime}}$ representing the attribution vectors explaining $\mathbf{x}$ and $\mathbf{x}^{\prime}$, and $f({\mathbf{x}})$ and $f({\mathbf{x}^{\prime}})$ the output prediction probabilities of $\mathbf{x}$ and $\mathbf{x}^{\prime}$, RIS and ROS metrics are defined as:
@@
$\forall\mathbf{x}^{\prime}\in\mathcal{N}_{\mathbf{x}}$, with ${p}$ defining the $l_{p}$ norm used to measure the changes and $\epsilon_{c}>0$ as a small values threshold to avoid zero division. The larger the RIS or ROS values, the more unstable the method is related to input or output prediction perturbation.
RIS and ROS handle values close to zero in normalization processes. However, the original versions implemented in Agarwal et al. [4] apply a clipping method that ignores negative values by trimming them to the small values threshold. Consequently, significant negative importance values are converted to a fixed threshold, leading the metrics to ignore instability in negative attributions. Then, we designed a clipping method that preserves significant values, either positive or negative, while discarding those close to zero.
@@
Furthermore, the original implementations of RIS and ROS may also ignore instances near the model’s decision boundaries. We solved that issue by sampling perturbations of each $\mathbf{x}$, sorting the perturbed data based on the distance to $\mathbf{x}$, and generating a neighborhood holding a minimum set of perturbed instances. This way, we ensure coverage and avoid measurement omissions or potential crashes in the metrics. Additionally, we extended RIS and ROS to evaluate mean changes in explanations.
To complement our stability evaluation, we introduce the Run Explanation Stability (RES) metric to assess the stability of local explainers for multiple runs on non-perturbed inputs. The principle behind RES is simple – it generates multiple explanations for the same original input under fixed model and explanation method settings and measures the extent to which such explanations change. According to the consistency property, identical inputs have identical explanations [42]. Given $n$ explanations $e_{\mathbf{x}i},\dots,e_{\mathbf{x}n}$ of a same non-perturbed instance $\mathbf{x}$, let $\bar{e}_{\mathbf{x}}$ be the mean explanation from those $n$ explanations, RES is defined as:
$\text{RES}(e_{\mathbf{x}},\bar{e}_{\mathbf{x}})=\max_{k=1}^{n}{\|\bar{e}_{\mathbf{x}}-e_{\mathbf{x}k}\|}\text{,}\>\>\>\forall\mathbf{x}\in\mathbf{X}\text{.}$ (13)
@@
Although we are introducing RES as part of the T-Explainer framework contributions, the rationale behind it was inspired by the Reiteration Similarity metric [7]; however, we simplified the formulation by using standard deviation as the similarity measure.
@@
RIS, ROS, and RES metrics are unitless quantifiers with no “ideal” desirable values. Reasonable values depend on the context of the application, meaning those metrics must be interpreted relatively by comparing their results across XAI methods [5]. However, lower values indicate methods with higher stability rates. Furthermore, RIS and ROS fit in the stability for slight variations strategy of the continuity category of XAI assessment, while RES fits in the consistency category [42].
A faithfulness metric evaluates the extent to which an explanation is faithful to the predictive model it explains. In this sense, we implemented the Prediction Gap on Important Features (PGI) metric to assess the methods’ faithfulness under the single/incremental deletion strategy for correctness [42]. Petsiuk et al. [44] introduced the intuition behind deletions, which Dai et al. [22] recently enhanced in the PGI metric. Given $top(k,e_{\mathbf{x}})$ being the $k$ most important features determined by a local explanation $e_{\mathbf{x}}$, we implemented PGI iterating on $m$ input instances as follows:
$\text{PGI}(\mathbf{x},f,e_{\mathbf{x}})=\frac{1}{m}\sum_{j=1}^{m}[1-|f(\mathbf{x})-f(\mathbf{\tilde{x}}_{j})|]$ (14)
where ${\tilde{x}_{i}}=0[i\notin top(k,e_{\mathbf{x}})]$ and $f$ returns predicted probabilities. PGI values will be in the $[0,1]$ range; the higher the value, the more faithful the explanations are compared to the “true” black-box behavior.
We also introduce the Local Accuracy Preservation (LAP) metric, which assesses the extent to which feature importance explainers preserve the local accuracy property, i.e., the model prediction should be reconstructed by the summation of the importance values (see Equation 5). Given an output prediction probability $f(\mathbf{x})$ and its respective explanation $g_{\mathbf{x}}$, we define the LAP metric as the frequency ratio in which $|f(\mathbf{x})-g_{\mathbf{x}}|\leq\epsilon$, $\forall\mathbf{x}\in\mathbf{X}$, with $\epsilon$ representing a tolerance error and $g$ following the additive modeling defined in Equation 4 [39]. The tolerance error is necessary since we do not expect any explanation modeling $g_{\mathbf{x}}$ to achieve a perfect match to the original model $f(\mathbf{x})$. LAP returns values in the $[0,1]$ range; the higher the value, the more faithful the explainer is regarding local accuracy preservation, i.e., the more an explainer keeps $g_{\mathbf{x}}$ close to $f(\mathbf{x})$, $\forall\mathbf{x}\in\mathbf{X}$, the more “locally accurate” it is.
Unlike previous implementations, we developed our metrics to run all the explainers under evaluation into the same execution cycles. It is an advantage regarding fair comparisons since it ensures the explainers are exposed to the same conditions. Our evaluations are based on five automated quantitative metrics, enabling users to compare T-Explainer’s performance with state-of-the-art feature importance methods formally. The metrics selected evaluate the consistency, continuity, and correctness dimensions of XAI evaluation [42], providing robust experimental results that go beyond the typical evaluations in XAI literature.